# Machine Learning Assisted Adjustment Boosts Inferential Efficiency of Randomized Controlled Trials

### v1

@@ Abstract
In this work, we proposed a novel inferential procedure assisted by machine learning based adjustment for randomized control trials. The method was developed under the Rosenbaum’s framework of exact tests in randomized experiments with covariate adjustments. Through extensive simulation experiments, we showed the proposed method can robustly control the type I error and can boost the inference efficiency for a randomized controlled trial (RCT). This advantage was further demonstrated in a real world example. The simplicity and robustness of the proposed method makes it a competitive candidate as a routine inference procedure for RCTs, especially when the number of baseline covariates is large, and when nonlinear association or interaction among covariates is expected. Its application may remarkably reduce the required sample size and cost of RCTs, such as phase III clinical trials.
@@ 1 Introduction
The randomized controlled trial (RCT) is the gold standard in determining treatment efficacy. Typically encompassing a substantial cohort size, particularly evident in phase III clinical trials, RCTs are characterized by their extensive participant involvement. However, this rigorous process comes at a considerable cost. Especially for trials on immunotherapy, the cost of some the newest therapies can reach \$850,000 per patient when including the value of medical support necessary to deliver these treatments [1]. Given the significant investment required, maximizing the statistical power of RCTs becomes critical.
@@
A common method used for baseline covariate adjustment is analysis of covariance (ANCOVA), which is essentially a linear model. In 2002, Rosenbaum proposed method of exact permutation inference which is free of distributional assumptions [2]. Although the method discussed was based on ordinary least square (OLS) linear regression, the author raised possibility of using more general forms of covariate adjustment. Nowadays, due to the ever decreasing cost of medical testing and sequencing techniques, the number of baseline data available in clinical trials has been growing dramatically. The rich data can provide a much finer picture of individuals and thus capture larger variation in the outcomes. This makes it possible to design inferential procedures with superior efficiency. However, the traditional adjustment methods using OLS method quickly break down when the number of covariates increases to a relatively large scale. Even if penalized regressions [3, 4, 5, 6] can acommodate high-dimensional data, such linear models are neither capable of handling nonlinear associations nor complex interactions.
@@
On the other hand, machine learning techniques have been successfully applied in numerous settings during the last decade. Supervised learning methods such as random forests (RFs), boosting machines, neural networks [7, 5, 8, 11] are known to be capable of handling high-dimensional data and model complex functional forms. In this work, we find the novel application of RF under the Rosenbaum’s framework is a valid and highly efficient adjustment method to boost the efficiency of inference for a RCT. The simplicity and robustness of the method makes it a potential routine inferential procedure for RCTs, especially when the number of baseline measurements is large, and when nonlinear association or between-covariate interaction is expected. The RF is used as a proof of concept, but it can be extended to more general modeling approches. Its application may remarkably reduce the required sample size and cost of RCTs, especially for phase III clinical trials.
@@ 2 Methods
In this work, we focus on the RCT with continuous outcomes and objective to compare two group means. Our method was based on the Rosenbaum’s framework. For details, please see Rosenbaum [2], and we here give a brief review. Suppose there are $n$ subjects, $j=1,\dots,n$, and the response of subject $j$ would be $r_{T_{j}}$ or $r_{C_{j}}$ if $j$ were assigned to treatment or control, respectively [9, 10]. The treatment effect can be written as $r_{T_{j}}-r_{C_{j}}$, and it can never be observed because subject $j$ can only receive one of the treatments. The treatment effect is additive if the treatment causes the response to change by a fixed amount $\tau$ for every $j$. The covariate vector $x_{j}$ describes the baseline characteristic of $j$. The variable $Z_{j}$ is the treatment indicator, where $Z_{j}=1,0$ indicates $j$ is assigned to the treatment or control group, respectively. Thus, the response of $j$ can be written as $R_{j}=Z_{j}r_{T_{j}}+(1-Z_{j})r_{C_{j}}$. Following Rosenbaum [2], we denote $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{n})^{T}$, $\mathbf{R}=(R_{1},R_{2},\dots,R_{n})^{T}$, $\mathbf{r}_{c}=(r_{C_{1}},r_{C_{2}},\dots,r_{C_{n}})^{T}$ and use $\mathbf{X}$ as the covariate matrix whose $j$th row is $\mathbf{x}_{j}^{T}$. The fixed but unobserved vector $\mathbf{r}_{C}$ are potential outcomes of all subjects if assigned to the controlled group, which and can be expressed as $\mathbf{r}_{C}=\mathbf{R}-\tau\mathbf{Z}$. Under this framework, the only stochastic component is the random assignment $\mathbf{Z}$. The test of $H_{0}:\tau=\tau_{0}$ can be performed using the statistic $t(\mathbf{Z},\mathbf{R}-\tau_{0}\mathbf{Z})$, where $\mathbf{R}-\tau_{0}\mathbf{Z}$ equals $\mathbf{r}_{C}$ under $H_{0}$ and additive treatment effects. Under this framework, the exact inference is the randomization inference derived from the randomization distribution of the statistics. The parametric distributions such as the normal distribution are used as approximations to randomization distributions [2].
@@
When covariate information $\mathbf{X}$ is available and correlate with the response variable $\mathbf{r}_{C}$, then we can write $\mathbf{r}_{C}=g(\mathbf{X})+\mathbf{e}$. The residual $\mathbf{e}$ is expected to have less variation than $\mathbf{r}_{c}$ if the baseline information explain part of the variation in the outcome. Note that if $H_{0}$ is true, we will have $\mathbf{r}_{C}=\mathbf{R}-\tau_{0}\mathbf{Z}$, and the residuals $\mathbf{e}$ can be used in place of $\mathbf{R}-\tau_{0}\mathbf{Z}$ for inference. Namely, the test statistic can be modified to $t(\mathbf{Z},\mathbf{e})$, which can lead to an improvement in efficiency due to smaller variation in $\mathbf{e}$.
Linear models are natural choice of $f$, but they tend to fail when the association is complex or the data is high dimensional. Here we investigate whether using modern machine learning method will achieve better efficiency for testing $H_{0}:\tau=\tau_{0}$ as compared to the straight two group comparison, while controlling the type I error rate properly. In this work, we select RF as the modeling method. For simplicity, let $\mathbf{Y}=(Y_{1},Y_{2},\dots Y_{3})^{T}$ denote the vector of continuous outcomes and note that we are interested in a regression task. The RF model is trained by repeatedly drawing bootstrap sample from the training data $(\mathbf{X},\mathbf{Y})$ for $B$ times. For the $b$th bootstrapped data $(\mathbf{X},\mathbf{Y})^{*}_{b}$ grow a random forest tree $T_{b}$ through recursive partitioning. A key feature of RF is that only $m<p$ variables will be randomly selected as candidates for each splitting, which reduces the correlations between trees and thereby improves the benefits of variance reduction by averaging. For a new observation $x$, the predicted response will be $\hat{f}(x)=\frac{1}{B}\sum_{b=1}^{B}T_{b}(x)$. For regression, the default $m=\lfloor\sqrt{p}\rfloor$, and the minimum node size is 5 [8].
@@
The RF approach was selected as a proof of concept due to its minimal requirement of parameter tuning and invariance to monotonic transformation of features [7]. More importantly, the out-of-bag (OOB) predictions can be used to attenuate the overfitting, thus avoiding manually splitting the data into training and validation set. Such simplicity makes it ideal for RCTs which typically requires detailed preplanning.
@@
    Use the data of all subjects and build a RF model using $\mathbf{X}$ as the features and $\mathbf{R-\tau_{0}\mathbf{Z}}$ as the response $\mathbf{Y}$.
@@
Under randomization and the assumption that the nonlinear models shift by some fixed amount $\bar{e}_{1}-\bar{e}_{0}=\bar{R}_{1}-\bar{R}_{0}$, so the average difference of the residuals by treatment group is the estimated treatment effect, while the residuals have a less dispersion. In machine learning, overfitting is a commonly encountered problem when a model has excessive flexibility and fits closely to the training data set. Such models are characterized with high variance and generalize poorly to new observations. This can also occur in modeling the relation between $\mathbf{X}$ and $\mathbf{R}-\tau_{0}\mathbf{Z}$ and lead to a biased estimation of the treatment effect.
@@ 3 Simulations
To investigate the performance of the proposed method, we conducted simulation studies under various scenarios for testing a one-sided hypothesis $H_{0}:\tau=0$ versus $H_{1}:\tau>0$. The results can be readily extended to two-sided tests. In each scenario, we generated a sample of $N=50,100,200$ and $400$ independent subjects, who are randomized into balanced two groups. There are $p=40$ mutually independent covariates following standard normal distributions. In each setting 10,000 Monte Carlo simulations were performed. The outcome $y$ was generated using the four different models:
@@
    $y=\tau Z+\beta[\sigma(x_{1}/2)-0.5]Z+\beta x_{2}+\beta x_{3}+\beta x_{4}+\epsilon,$
@@
where $Z$ is the treatment indicator, $\tau$ is the treatment effect, $\sigma(x)=\exp(x)(1+\exp(x))^{-1}$ is the sigmoid function, $x_{k}$ is the $k$th covariate, and $\epsilon$ is the error term. The subscript for subject is omitted. We examined $\beta=0.2,0.5,0.8$ to represent different association strength between covariates and the outcome, as well as error terms from standard normal, log-normal, and Gumbel distributions. The straight two-sample $t$-test and Wilcoxon rank-sum test are commonly used analytical methods. We also included $t$-test and Wilcoxon rank-sum test on the residuals from linear regression on all baseline covariates ($t$-test-LM and Wilcoxon-LM). Finally, we included the proposed RF adjusted methods, which use the $t$-test and Wilcoxon rank-rum test on the residuals from RF model ($t$-test-RF and Wilcoxon-RF). We examined the type I error control with $\tau=0$, and the statistical power with $\tau=0.3$ and $0.6$.
@@
Figures 1$\sim$4 show the simulation results from models with normal error terms. It is shows that all tests robustly control the type I error rates when $\tau=0$ in all settings. In models with primary nonlinear associations (Model 1, 2, 4), the RF adjusted tests ($t$-test-RF and Wilcoxon-RF) show uniformly higher power than the other models. In Figure 1, bottom right panel, the RF tests attains over 90% power at around $N=200$, while the other test would require at least 100 more subjects to achieve the same power. In models with primary linear associations (Figure 3, Model 3), the performance of RF tests is mostly comparable with that of the linear model adjusted models. Only when the association is very strong (Beta=0.8), the linear model methods will have a advantage of around 7%. It should be noted that under some scenarios the linear model tend to hurt the efficiency, making its power lower than the straight $t$-test and Wilcoxon test, even when the true association is primarily linear. On the other hand, the RF tests consistently have similar or better performance versus the unadjusted tests. This suggests the application of RF adjustment is generally safe and unlikely to result in a loss of efficiency. Figure 5 shows that when the distribution of error is non-normal, all tests still control the type I error robustly. Meanwhile, the RF tests still exhibit similar advantage over other methods, but the Wilcoxon-RF shows a further improvement over the $t$-test-RF under non-normality. The results on cases with log-normal errors are similar (Supplement). In addition, although the Rosenbaum’s framework assumes an additive treatment effect, the proposed method show equally well type I error control and efficiency improvement when there are interactions between treatment and covariates (Figure 4). The results are similar is the randomization is unbalanced (data not shown). Overall, our simulations suggest Wilcoxon-RF test as the best option among the methods tested.
@@
Figure 1: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 1 and with error terms from standard normal distributions.
Figure 2: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 2 and with error terms from standard normal distributions.
Figure 3: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 3 and with error terms from standard normal distributions.
Figure 4: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 4 and with error terms from standard normal distributions.
@@
Figure 5: Type I error (Tau=0) and statistical power of the testing procedures under different settings of Model 4 and with error terms from standard Gumbel distributions.

### v2

# Machine Learning Assisted Adjustment Boosts Efficiency of Exact Inference in Randomized Controlled Trials

@@ Abstract
In this work, we proposed a novel inferential procedure assisted by machine learning based adjustment for randomized control trials. The method was developed under the Rosenbaum’s framework of exact tests in randomized experiments with covariate adjustments. Through extensive simulation experiments, we showed the proposed method can robustly control the type I error and can boost the statistical efficiency for a randomized controlled trial (RCT). This advantage was further demonstrated in a real-world example. The simplicity, flexibility, and robustness of the proposed method makes it a competitive candidate as a routine inference procedure for RCTs, especially when nonlinear association or interaction among covariates is expected. Its application may remarkably reduce the required sample size and cost of RCTs, such as phase III clinical trials.
@@ 1 Introduction
The randomized controlled trial (RCT) is the gold standard in determining treatment efficacy. RCTs are characterized by their extensive participant involvement, which typically encompass a substantial cohort size, particularly evident in phase III clinical trials. However, this rigorous process comes at a considerable cost. Especially for trials on immunotherapy, the cost of some of the newest therapies can reach \$850,000 per patient when including the value of medical support necessary to deliver these treatments [1]. Given the significant investment required, maximizing the statistical power of RCTs becomes critical.
A common approach to improve the statistical efficiency is baseline adjustment using linear models, such as analysis of covariance (ANCOVA). Related models and their applications in RCT has been extensively discussed [2, 3]. The method of covariate adjustments is supported by FDA in phase III oncology trials due to the enhanced efficiency and its minimal impact on bias or type I error rate [5]. Hence, the agency advises sponsors to incorporate adjustments for covariates expected to show the strong association with the outcome of interest. In 2002, Rosenbaum proposed method of exact inference which is free of distributional assumptions [4]. This method is endorsed by the FDA guidance, which states that ”Sponsors can conduct randomization/permutation tests with covariate adjustment (Rosenbaum 2002)”. Although this method was based on ordinary least square (OLS) linear regression, the author raised possibility of using more general forms of covariate adjustment.
@@
Nowadays, due to the ever-decreasing cost of medical testing and sequencing techniques, the number of baseline data available in clinical trials has been growing dramatically. The rich data can provide a much finer picture of individuals and thus capture larger variation in the outcomes. This makes it possible to design inferential procedures with superior efficiency. However, the traditional adjustment methods using OLS method quickly break down when the number of covariates increases to a relatively large scale. Even if penalized regressions [6, 7, 8, 9] can accommodate high-dimensional data, such linear models are neither capable of handling nonlinear associations nor complex interactions.
@@
On the other hand, machine learning techniques have been successfully applied in numerous settings during the last decade. Supervised learning methods such as random forests (RFs), boosting machines, neural networks [10, 8, 11, 12] are known to be capable of handling high-dimensional data and model complex functional forms. While efforts have been made towards the application of machine learning models for the covariates adjustment in randomized studies, most of the works focused on estimations and asymptotics [13, 14, 15]. In this work, we propose a novel approach of machine learning based covariate adjustment under the Rosenbaum’s framework, which makes exact inference possible [4]. The proposed method focuses on continuous outcomes, and is shown to be an unbiased and flexible adjustment method to boost the statistical efficiency for a RCT. These characteristics make it a novel and powerful inferential procedure for RCTs, especially when nonlinear association or between-covariate interaction is expected. The RF model is used under this framework as a proof of concept, but it can be extended to more general modeling approaches. Its application may remarkably reduce the required sample size and cost of RCTs, especially for phase III clinical trials.
@@ 2.1 Hypothesis testing under non-parametric adjustment
In this work, we focus on RCTs with continuous outcomes and with an objective to compare two group means. Our method is based on the Rosenbaum’s framework. For details, please see Rosenbaum [4]. As a brief background, suppose there are $n$ subjects, $j=1,\dots,n$, and the response of subject $j$ would be $r_{T_{j}}$ or $r_{C_{j}}$ if $j$ were assigned to treatment or control, respectively [16, 17]. The treatment effect can be written as $\tau_{i}=r_{T_{j}}-r_{C_{j}}$, and it can never be observed because subject $j$ can only receive one of the treatments. The treatment effect is additive if the treatment causes the response to change by a fixed amount $\tau$ for every $j$. The covariate vector $x_{j}$ describes the baseline characteristic of $j$. The variable $Z_{j}$ is the treatment indicator, where $Z_{j}=1,0$ indicates $j$ is assigned to the treatment or control group, respectively. Thus, the response of $j$ can be written as $R_{j}=Z_{j}r_{T_{j}}+(1-Z_{j})r_{C_{j}}$. Following Rosenbaum [4], we denote $\mathbf{Z}=(Z_{1},Z_{2},\dots,Z_{n})^{T}$, $\mathbf{R}=(R_{1},R_{2},\dots,R_{n})^{T}$, $\mathbf{r}_{c}=(r_{C_{1}},r_{C_{2}},\dots,r_{C_{n}})^{T}$ and use $\mathbf{X}$ as the covariate matrix whose $j$th row is $\mathbf{x}_{j}^{T}$. The fixed but unobserved vector $\mathbf{r}_{C}$ are potential outcomes of all subjects if assigned to the controlled group, which can be expressed as $\mathbf{r}_{C}=\mathbf{R}-\tau\mathbf{Z}$. Under this framework, the only stochastic component is the random assignment $\mathbf{Z}$. The test of $H_{0}:\tau=\tau_{0}$ can be performed using the statistic $t(\mathbf{Z},\mathbf{Y})$, where $\mathbf{Y}=\mathbf{R}-\tau_{0}\mathbf{Z}$. The response variable $\mathbf{Y}$ is the observed outcome $\mathbf{R}$ adjusted with respect to $\tau_{0}$, which will be the variable used in the inference procedure. Thus we have $\mathbf{Y}=\mathbf{R}$ when $\tau_{0}=0$. Under this framework, the exact inference is the randomization inference derived from the randomization distribution of the statistics. The parametric distributions such as the normal distribution are used as approximations to randomization distributions [4].
@@
When covariate information $\mathbf{X}$ is available and associated with the response variable $\mathbf{Y}$, then we can write $\mathbf{Y}=g(\mathbf{X})+\mathbf{e}$. The residual $\mathbf{e}$ is expected to have less variation than $\mathbf{Y}$ if the baseline covariates explain part of the variation, and it can be used in place of $\mathbf{Y}$ for inference. Thus the test statistic can be modified to $t(\mathbf{Z},\mathbf{e})$, which can lead to an improvement in efficiency due to smaller variation in $\mathbf{e}$.
Linear models are natural choice of $g$, but they tend to fail when the association is complex or the data is high dimensional. Here we investigate whether using modern machine learning method will achieve better efficiency for testing $H_{0}:\tau=\tau_{0}$ as compared to the straight two group comparison, while controlling the type I error rate properly. In this work, we select RF as the exact inference approach.
@@
The RF approach was selected for our approach due to its minimal requirement of parameter tuning and invariance to monotonic transformation of features [10]. More importantly, the out-of-bag (OOB) predictions can be used to attenuate the overfitting, thus avoiding manually splitting the data into training and validation set. Such simplicity makes it ideal for RCTs, which typically requires detailed preplanning.
@@
    Use the data from all subjects and build a RF model using $\mathbf{X}$ as the features and $\mathbf{R-\tau_{0}\mathbf{Z}}$ as the response $\mathbf{Y}$.
@@
2.2 Estimation
Without loss of generality, we focus on the null hypothesis $H_{0}:\tau_{0}=0$ and relax the additivity assumption for the discussion of estimation. When $H_{0}:\tau_{0}=0$, we have $\mathbf{Y}=\mathbf{R}$ which is the observed response. Conceptually, the adjustment by $g(\mathbf{X})$ shifts the response at any point in the covariate space by the same amount for both arms. Therefore, the difference in the residuals should be the same as that in the original response, so the treatment effect is estimated as,
$\hat{\tau}$ $=\frac{1}{n_{1}}\sum_{Z_{i}=1}e_{i}-\frac{1}{n_{0}}\sum_{Z_{i}=0}e_{i}=\frac{1}{n_{1}}\sum_{i=1}^{n}(Y_{i}-g(X_{i}))Z_{i}-\frac{1}{n_{0}}\sum_{i=i}^{n}(Y_{i}-g(X_{i}))(1-Z_{i}),$ (2.1)
where $n_{0}$ and $n_{1}$ are numbers of subjects in control and treatment groups. The same estimator has been proposed by Opper [14] for a different inference procedure. The condition for $\hat{\tau}$ in Equation 2.1 to be unbiased estimator was also given [14]. In our framework, the estimation of $g(\mathbf{X})$ does not involve treatment assignment, so it is invariant under randomization. Therefore, the only stochastic component is $Z_{i}$. Therefore, we can write $Y_{i}=r_{C_{i}}+Z_{i}\tau_{i},i=1,2,\dots n$, then the expectation of estimated treatment effect is given as,
$E(\hat{\tau})$ $=E(\frac{1}{n_{1}}\sum_{i=1}^{n}(Y_{i}-g(X_{i}))Z_{i}-\frac{1}{n_{0}}\sum_{i=i}^{n}(Y_{i}-g(X_{i}))(1-Z_{i}))$ (2.2)
$=\frac{1}{n}\sum_{i=1}^{n}\tau_{i}-\frac{1}{n}\sum_{i=1}^{n}(E(g(X_{i})|_{Z_{i}=1})-E(g(X_{i})|_{Z_{i}=0})).$
Under additivity, we have $\tau_{i}=\tau,i=1,2,\dots,n$. Thus, $E(\hat{\tau})=\tau+\frac{1}{n}\sum_{i=1}^{n}(g(X_{i})|_{Z_{i}=1}-g(X_{i})|_{Z_{i}=0})$. Therefore, a sufficient condition for $\hat{\tau}$ to be an unbiased estimator, i.e., $E(\hat{\tau})=\tau$ is
$E(g(X_{i})|_{Z_{i}=0})=E(g(X_{i})|_{Z_{i}=1}).$
One would assume that this condition is natually fulfilled since the estimation of $g$ does not involve treatment assignment. However, it can be violated particularly if $g$ overfits the data. Overfitting occurs when $g(X_{i})$ is overly close to the observed response $Y_{i}$, meaning that as the flexibility of $g$ increases, we have $g(X_{i})\to r_{C_{i}}$ when $Z_{i}=0$, and $g(X_{i})\to r_{C_{i}}+\tau_{i}$ when $Z_{i}=0$. Consequently, we have $E(g(X_{i})|_{Z_{i}=1})-E(g(X_{i})|_{Z_{i}=0})\to\tau_{i}$ and $E(\hat{\tau})\to 0$. This suggests when $g$ overfits the data, the estimated treatment effect will be biased towards 00. As long as the machine learning model is adequately flexible, it can always overfit the data even if the sample size is large (e.g., the $k$ Nearest Neighbor with $k=1$). Therefore, for models with high complexity, the bias term is $O(1)$ as $n\to\infty$, thus simply increasing the sample size does not necessarily mitigate the bias. One way to counter this problem is to use the predictions from models not trained on the observations at question, which can be easily achieved through cross-validations, or OOB predictions.
Next, we derive the variance of the permutation distribution of $\hat{\tau}$. Note that $\hat{\tau}$ can be rewritten as,
$\hat{\tau}$ $=\sum^{n}_{i=1}\frac{ne_{i}Z_{i}}{n_{1}n_{0}}-\frac{1}{n}\sum^{n}_{i=1}e_{i}.$ (2.3)
Recall that in our method, the residuals are fixed after the estimation of $g$. Therefore, the variance of the estimator is given as,
$Var(\hat{\tau})$ $=\frac{1}{n_{1}n_{0}}\sum^{n}_{i=1}e_{i}^{2}.$ (2.4)
The equation above suggests the variance will be smallest when $\sum^{n}_{i=1}e_{i}^{2}$ is minimized, which corresponds to a model $g$ that is optimized towards minimal mean squared error. This observation bridges the inference problem with classic predictive modeling problems. It is notable that Opper obtained the same form of variance in his work[14]. However, the author ignored the variability of $g(\mathbf{X})$, and assumed the residuals to be fixed, which does not align with their procedure that requires estimation of $g(\mathbf{X})$ under each randomization. Interestingly, the same expression is correct under our framework, where the residuals are indeed fixed because estimation of $g(\mathbf{X})$ does not depend on $\mathbf{Z}$.
When the treatment effect is non-additive, the treatment effect is defined as $\tau=\frac{1}{n}\sum_{i=1}^{n}\tau_{i}$. Therefore, the above results on unbiasedness still hold. On the other hand, non-additivity implies that there are interactions between treatment and covariates. Thus the function $g(\mathbf{X})$ can be better described as two separate functions for two arms, namely $g_{0}(\mathbf{X})$ and $g_{1}(\mathbf{X})$. Although testing the null hypothesis under the non-additivity assumption is not very well defined, a natural question under non-additivity is whether modeling $g_{0}(\mathbf{X})$ and $g_{1}(\mathbf{X})$ individually will gain efficiency over estimating a common regression function $g(\mathbf{X})$ by pooling the observations from two arms together. Such a strategy was adopted by Wager[13] and Opper[14]. To investigate this problem, we examined the estimator of treatment effect from the cross-estimation (CE) method by Wager [13],
$\hat{\tau}_{CE}=\frac{1}{n}\sum_{i=1}^{n}(g_{1}(X_{i})-g_{0}(X_{i}))+\sum_{Z_{i}=1}\frac{Y_{i}-g_{1}(X_{i})}{n_{1}}-\sum_{Z_{i}=0}\frac{Y_{i}-g_{0}(X_{i})}{n_{0}}.$ (2.5)
Here, we ignored the $(-i)$ notation, which was used to denote that the estimator of $g$ was not dependent on the $i$th training ovservation. Then based on their result, the variance of the estimator is given as,
$Var(\hat{\tau}_{CE})$ $=\sum_{z\in(0,1)}\sum_{Z_{i}=z}\frac{(Y_{i}-(1-p)g_{1}(X_{i})-pg_{0}(X_{i}))^{2}}{n_{z}(n_{z}-1)},$ (2.6)
where $n_{z}={n\choose n_{0}}$ is the number of all unique randomizations and $p=n_{1}/n$. Note that the variance depends on $g_{0}$ and $g_{1}$ only through a weighted sum $g(X_{i})=(1-p)g_{0}(X_{i})+pg_{1}(X_{i})$, which can be estimated through pooling the observations from two arms and proper weighting. Therefore, even under non-additivity, the knowledge of $g_{0}$ and $g_{1}$ does not lead to a further gain in efficiency, so the estimation of $g(\mathbf{X})$ across two components is unnecessary. In fact, under a balanced design, the sample size for training the model will be halved to train the models for the control and treatment groups separately, which can result in reduced model performance and consequently lower inference efficiency. Therefore, the proposed method is not only more straightforward for exact inference, but it also achieves a better efficiency through pooling observations from both intervention arms.
@@ 3 Simulations
To demonstrate the performance of the proposed method, we conducted a set of comprehensive simulation studies under various scenarios for testing a one-sided hypothesis $H_{0}:\tau=0$ versus $H_{1}:\tau>0$. The results can be readily extended to two-sided tests. In each scenario, we generated samples of size $N=50,100,200$ and $400$ independent subjects, who were then randomized into two balanced groups. There were $p=40$ mutually independent covariates following standard normal distributions. In each setting 10,000 Monte Carlo simulations were performed. The outcome $y$ was generated using the four different models:
@@
    $y=\tau Z+\beta[\sigma(x_{1}/2)-1/2]Z+\beta x^{2}_{2}+\beta cos(x_{3})+\beta x_{4}+\epsilon,$
@@
where $Z$ is the treatment indicator, $\tau$ is the treatment effect, $\sigma(x)=\exp(x)(1+\exp(x))^{-1}$ is the sigmoid function, $x_{k}$ is the $k$th covariate, and $\epsilon$ is the error term. The subscript for subject is omitted. We examined $\beta=0.2,0.5,0.8$ to represent different levels of association between the covariates and the outcome measure. In addition, we utilized error terms from standard normal, log-normal, and Gumbel distributions. For comparing the outcomes between the two treatment ars we utilized the Wilcoxon rank-sum test and the two-sample $t$-test. We also included $t$-test and Wilcoxon rank-sum test on the residuals from linear regression on all baseline covariates ($t$-test-LM and Wilcoxon-LM). Finally, we included the proposed RF adjusted methods, which use the $t$-test and Wilcoxon rank-rum test on the residuals from RF model ($t$-test-RF and Wilcoxon-RF). The Wager’s cross-estimation (CE) method was also included for comparison, where a one-sided $Z$-test was constructed based on the estimate and standard error from the ate.randomForest function in R crossEstimation package [13]. We examined the type I error control with $\tau=0$, and the statistical power with $\tau=0.3$ and $0.6$. The results of all simulations are shown in the supplementary materials.
@@
Figures 1 shows that all tests robustly control the type I error rates when $\tau=0$ under non-Gaussian error terms. In models with primary nonlinear associations, the RF adjusted tests ($t$-test-RF and Wilcoxon-RF) show higher power than the other methods. In the right panel of Figure 2, we see that the RF tests attains over 80% power at around $N=170$, while the other tests require approximately 80 more subjects to achieve the same power. In models with primary linear associations, the performance of RF tests is mostly comparable with that of the linear adjusted methods. Only when the association is very strong ($\beta=0.8$), the linear model methods will have a advantage of around 7% (Figures S3 & S7). It should be noted that under some scenarios (e.g. small sample sizes), the linear model approach tends to hurt the efficiency, having power lower than the straight $t$-test and Wilcoxon test, even when the true association is linear. On the other hand, the RF tests have similar or better performance than the unadjusted tests in most cases. One exception is the setting with weak associations and log-normal errors (Figure 3). However, as the association becomes stronger, the RF test still outperforms other methods. In addition, although the Rosenbaum’s framework assumes an additive treatment effect, the proposed method shows equally well type I error control and efficiency improvement when there are interactions between treatment and covariates (Figures S4, S8, S12). It is also notable that our proposed method consistently outperforms the CE method [13], even when the covariate by treatment interaction is present. This demonstrates that the explicit modeling of $g$ separately for two arms using nonparametric methods reduces statistical efficiency due to the smaller training sample sizes. The results are similar when the randomization is unbalanced (data not shown). Overall, our simulations suggest that, in most cases, the Wilcoxon-RF test as the best option among the methods tested.
@@
Figure 1: Type I error of the testing procedures under Model 1 with $\tau=0,\beta=0.8$ and error terms from Gumbel and log-normal distributions. Dashed line indicates level of 0.05.
@@
Figure 2: The statistical power of the testing procedures under different settings of Model 1 and with error terms from Gumbel distributions. Dashed line indicates power of 0.8.
Figure 3: The statistical power of the testing procedures under different settings of Model 1 and with error terms from log-normal distributions. Dashed line indicates power of 0.8.